Solve g(y)= 1/y over the interval (1,4)

1
Expert's answer
2022-01-12T17:20:29-0500

Average rate of change= "\frac{g(b)-g(a)}{b-a}\\"

"=\frac{g(4)-g(1)}{4-1}\\
=\frac{\frac{1}{4}-1}{3.}\\
=\frac{-1}{4}."